OpenStudy (anonymous):

Why is m1=3, m2=-1/3 a perpendicular line, but m1=5.5/2.7, m2=(2.7/5.5), with a negative 1 exponent outside (2.7/5.5), a parrallel line?

OpenStudy (anonymous):

Parallel lines have the same slope. Perpendicular lines have slopes which are negative reciprocals of each other.

OpenStudy (anonymous):

Both appeared to be perpendicular to me...

OpenStudy (anonymous):

No because \[m2=(\frac{2.7}{5.5})^{-1} = \frac{2.7^{-1}}{5.5^{-1}} = \frac{5.5}{2.7} = m1\]

OpenStudy (anonymous):

A negative 1 exponent is not the same as multiplying by -1.

OpenStudy (anonymous):

Right, I know that..

OpenStudy (anonymous):

Ok, thanks anyway....hopefully I'll be able to finish this problem set ok now. :)

OpenStudy (anonymous):

Does what I said make sense?

OpenStudy (anonymous):

If you simplify m2 you get the same value as m1. Since the slopes are the same, the lines are parallel.